Hahn Company uses a job order costing system. Its ide predetermined overhead rate uses direct labor-hours as the alloc base. The company pays its direct laborers $20.00 per hour. During the year, the company started and completed only two jobs-Job Alpha, which used 66,700 direct labor-hours, and Job Omega. The job cost sheets for the these two jobs are shown below: Job Alpha Direct materials Direct labor Manufacturing overhead applied 2,348,000 Total job cost Job omega 455,000 Direct materials 600,000 Direct labor 390,000 Manufacturing overhead applied 1,445,000 Total job cost. Required 1. Calculate the plantwide predetermined overhead rate. 2. Complete the job cost sheet for Job Alpha.
Answer #1
1
Direct labor hours used by Job-Omega = 600000/20= 30000
Plantwide predetermined overhead rate = manufacturing overhead applied/Direct labor hours
=390000/30000= 13 per DLH
2
Job Alpha
Direct materials 146900
Direct labor 1334000 =66700*20
Manufacturing overhead applied 867100 =66700*13
Total job cost 2348000